OpenStudy (anonymous):

A rectangle is 5x as long as it is wide. If it were 24 cm shorter and 24 cm wider, it would be square. what are it's dimentions?

OpenStudy (anonymous):

dimentions*`

OpenStudy (anonymous):

OMG Dimensions

OpenStudy (anonymous):

I know l=lent and w=w

OpenStudy (anonymous):

length, gosh what is wrong with me today?

OpenStudy (anonymous):

L=5W L-24=W+24 5W-24=W+24 5W-W=24+24 4W=48 W=12 L=60 CHECKING: 60-24=12+24 36=36 CHECKED

OpenStudy (anonymous):

i see, tyvm!(: Here's a medal x]

OpenStudy (anonymous):

TNX A LOT

OpenStudy (anonymous):

sure? xD

OpenStudy (anonymous):

SURE ON WHAT?